From a standard deck of 52 cards, 5 cards are dealt and the order of the cards is unimportant. In how many ways can you receive all face cards?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

From a standard deck of 52 cards, 5 cards are dealt and the order of the cards is unimportant. In how many ways can you receive all face cards?
-----
There are 40 face cards:
Ans: 40C5 = (40*39*38*37*36)/(1*2*3*4*5) = 658,008
===========================
Cheers,
Stan H.

Answer by Alan3354(69443)

From a standard deck of 52 cards, 5 cards are dealt and the order of the cards is unimportant. In how many ways can you receive all face cards?
-----------------
There are 12 face cards, J, Q and K of each suit.
12C5 = 12*11*10*9*8/5*4*3*2*1
= 792